If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
Answer by stanbon(48523)

If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
-----------------------------
Sam rate = 1/4 job/day
Lisa rate = 1/6 job/day
Tome rate = 1/2 job/day
----------------------------
Let together rate be 1/x job/day
-----------------------------
Equation:
rate + rate + rate = together rate
1/4 + 1/6 + 1/2 = 1/x
(3 + 2 + 6)/12 = 1/x
11/12 = 1/x
11x = 12
x = 12/11 = 1.11 days to do the job together.
===============================
Cheers,
Stan H.